# Anticipating the Selectivity of Cyclization Reaction Pathways with Neural Network Potentials

### v1

@@ 3.3 Retrospective analysis of key steps in natural product synthesis planning
In natural product synthesis, there is often an enabling reaction that enables the construction of the molecular scaffold along the synthesis route; for a subset of products, this may be a cyclization that simultaneously sets multiple stereocenters.79 In recent years, detailed computational analysis of proposed synthetic steps have been used to explain selectivity and guide synthetic efforts 7, 8, 9, 10, 11, 12, 13, thus we see an opportunity to apply RAMP to the assessment of feasibility of these enabling reactions during natural product route planning as a way to derisk experiments and reveal potential unwanted byproducts.
@@ 3.3.1 Synthesis of a key salvinorin intermediate
As our validations thus far have demonstrated that RAMP is particularly compatible with the evaluation of Diels-Alder cycloadditions, the first synthesis planning we explored was a key step of Zimdars’ synthesis of salvinorin A 80 (Figure 4A). In this synthesis, a single concerted step enables the generation of two cycles and two stereocenters. The preferred pathway found in our analysis is the intramolecular Diels-Alder reaction of (S,E)-1-(furan-3-yl)-3,5-dimethylhexa-3,5-dien-1-yl acrylate (15) to a lactone (16) (Figure 4C) which matches the experimental results from Zimdars et al. 80.
RAMP identifies two other diastereomers (Figure S10) of the same lactone with one of the diastereomers (17) having a barrier only 1.6 kcal/mol greater than the product (16) which would indicate this diastereomer would appear as a minor product. However, this step is experimentally reported to be completely selective for 16 indicating minor disagreement between experiment and our calculated results. Despite this, the major product for this reaction is still identified as the preferred pathway in our analysis.
@@ 3.3.2 Synthesis of endiandric acid C
Key steps that span multiple elementary steps, as in cascade reactions, may be more challenging to predict a priori through intuition alone. We performed a mechanistic search starting from the complex polyene 18 from the last step in the Nicolaou’s synthesis of endiandric acid C, which involves three sequential cyclization reactions 81. These elementary steps comprise a b4f4 cyclization as well as a stereodetermining step that yields a bridged ring.
From the reactant 18, the first step that is found is an 8 contrarotary cyclization to form an 8 membered ring (19). This is followed by a 6 disrotary cyclization. Finally, a Diels-Alder reaction is used to form the bridged ring that is in the product (21). With the stereoisomers enumerated from RAMP, we are able to successfully determine the correct stereoisomer for each of these steps using the calculated barriers (Figure 5C). There is quantitative disagreement between the barrier heights estimated by AIMNet2-rxn and those estimated using its reference method, justifying our inclusion of DFT as the final evaluation during each single-step exploration; importantly, the ranking of outcomes remains consistent.
Intermediates after filtering | Wall time (h) | ||||||||
Case Study | N | bnfn | Total | Stability | Pre-DFT | DFT | |||
Benchmark - 1 | 20 | 3 | 14120 | 8714 | 3515 | 750 | 545 | 5.1 | 21.3 |
Benchmark - 2 | 24 | 4 | 4030 | 1995 | 1880 | 1206 | 198 | 0.7 | 14.6 |
DA - 1 | 33 | 3 | 49548 | 38006 | 14357 | 2217 | 531 | 6.0 | 36.9 |
DA - 2 | 34 | 3 | 80290 | 53683 | 5812 | 514 | 173 | 2.1 | 57.7 |
Salvinorin A | 36 | 3 | 23582 | 15548 | 12571 | 5281 | 690 | 13.1 | 23.7 |
Endiandric acid C | 52 | 4 | 911946 | 775230 | 521271 | 85068 | 402 | 62.8 | 298.7 |
@@ 4 Conclusion
In this article, we have demonstrated the productive combination of graph-based enumeration and neural network potentials for the computational evaluation of complex cyclization reactions with relevance to natural product synthesis planning. We implement our workflow in an open-source tool, RAMP, which by default performs a “b4f4” and stereoenumeration scheme to generate a broad range of potential outcomes. To cope with the large number of possible elementary steps, we employ inexpensive machine learning models for intermediate filtering and perform subsequent thermodynamic and kinetic assessment with the reactive NNP AIMNet2-rxn. Many elements of this workflow take direct inspiration from YARP. We illustrate the successful construction of complex cyclization pathways with various case studies: a stereoselectivity study of an intramolecular Diels-Alder reaction and two examples from natural product syntheses representing various levels of elementary step complexity.
The summary statistics presented in Table 1 demonstrate RAMP’s ability to explore these mechanisms with manageable computational cost. One clear trend that emerges is that the number of enumerated intermediates is very sensitive to the number of atoms which compounds the existing challenges associated with modeling larger reacting systems. RAMP partially mitigates this combinatorial explosion with its progressive filtering stages which help reduce the number of DFT transition state optimizations by at least an order of magnitude in all case studies. Despite the efficacy of these filters, the wall time associated with performing the DFT calculations is still quite high in cases with larger molecules, highlighting the challenges associated with modeling systems of this size and complexity.
The rise of NNPs capable of providing DFT-quality energies for equilibrium and transition state geometries will continue to assist mechanistic modeling. At present, the strength of RAMP is in the analysis of neutral, closed-shell reactions, particularly pericyclic reactions like electrocyclizations and cycloadditions. Currently, RAMP is incapable of handling elements outside of C, H, O, and N, charged species, and open-shelled chemistry. This represents a significant limitation of its utility beyond analyzing the particular reaction types in this study. However, due to the modular nature of RAMP, it will be straightforward to exchange the current NNP for those that will be released in coming years that will be capable of handling a larger scope of reactions.
@@
The code and data that support the findings of this study are openly available at https://github.com/ncasetti7/RAMP/tree/main. The neural network potential used in this study will be available when its corresponding paper 41 is published.

### v2

@@ 3.3 Retrospective analysis of key steps in natural product synthesis planning
In natural product synthesis, there is often an enabling reaction that enables the construction of the molecular scaffold along the synthesis route; for a subset of products, this may be a cyclization that simultaneously sets multiple stereocenters.81 In recent years, detailed computational analysis of proposed synthetic steps have been used to explain selectivity and guide synthetic efforts 7, 8, 9, 10, 11, 12, 13, thus we see an opportunity to apply REVAMP to the assessment of feasibility of these enabling reactions during natural product route planning as a way to derisk experiments and reveal potential unwanted byproducts.
@@ 3.3.1 Synthesis of a key salvinorin intermediate
As our validations thus far have demonstrated that REVAMP is particularly compatible with the evaluation of Diels-Alder cycloadditions, the first synthesis planning we explored was a key step of Zimdars’ synthesis of salvinorin A 82 (Figure 4A). In this synthesis, a single concerted step enables the generation of two cycles and two stereocenters. The preferred pathway found in our analysis is the intramolecular Diels-Alder reaction of (S,E)-1-(furan-3-yl)-3,5-dimethylhexa-3,5-dien-1-yl acrylate (15) to a lactone (16) (Figure 4C) which matches the experimental results from Zimdars et al. 82.
REVAMP identifies two other diastereomers (Figure S11) of the same lactone with one of the diastereomers (17) having a barrier only 1.6 kcal/mol greater than the product (16) which would indicate this diastereomer would appear as a minor product. However, this step is experimentally reported to be completely selective for 16 indicating minor disagreement between experiment and our calculated results. Despite this, the major product for this reaction is still identified as the preferred pathway in our analysis.
@@ 3.3.2 Synthesis of endiandric acid C
Key steps that span multiple elementary steps, as in cascade reactions, may be more challenging to predict a priori through intuition alone. We performed a mechanistic search starting from the complex polyene 18 from the last step in the Nicolaou’s synthesis of endiandric acid C, which involves three sequential cyclization reactions 83. These elementary steps comprise a b4f4 cyclization as well as a stereodetermining step that yields a bridged ring.
From the reactant 18, the first step that is found is an 8 contrarotary cyclization to form an 8 membered ring (19). This is followed by a 6 disrotary cyclization. Finally, a Diels-Alder reaction is used to form the bridged ring that is in the product (21). With the stereoisomers enumerated from REVAMP, we are able to successfully determine the correct stereoisomer for each of these steps using the calculated barriers (Figure 5C). The accuracy of AIMNet2-rxn compared to the reference DFT can vary depending on the chemical space and reaction mechanism coverage. We therefore use DFT as the final evaluation during each single-step exploration to provide a consistent physics-based evaluation; importantly, the ranking of outcomes remains consistent, empirically. Another justification for the inclusion of DFT as a final evaluation are the N/A barrier values in Figure 5. These indicate that the DFT transition state optimization did not converge to a valid transition state which means that either the AIMNet2-rxn transition state conformer was too far away from the corresponding DFT transition state conformer or that the corresponding DFT transition state conformer doesn’t exist. In the former case, increasing the number of transition states sampled by AIMNet2-rxn or incorporating global optimization techniques (like Monte Carlo multiple minimum sampling 84) would help increase the reliability of DFT transition state convergence at an increased computational expense. In the latter case, we assume that the DFT PES is a better representation of the ground truth PES thus allowing us to label these transition states and their corresponding elementary steps as kinetically irrelevant.
| Intermediates after filtering | Wall time (h) | |||||||||
| Case Study | N | bnfn | Total | Stability | Pre-DFT | DFT | ||||
| Benchmark - 1 | 20 | 3 | 14120 | 8714 | 3515 | 750 | 545 | 97 | 5.1 | 21.3 |
| Benchmark - 2 | 24 | 4 | 4030 | 1995 | 1880 | 1206 | 198 | 60 | 0.7 | 14.6 |
| DA - 1 | 33 | 3 | 49548 | 38006 | 14357 | 2217 | 531 | 22 | 6.0 | 36.9 |
| DA - 2 | 34 | 3 | 80290 | 53683 | 5812 | 514 | 173 | 20 | 2.1 | 57.7 |
| Salvinorin A | 36 | 3 | 23582 | 15548 | 12571 | 5281 | 690 | 28 | 13.1 | 23.7 |
| Endiandric acid C | 52 | 4 | 911946 | 775230 | 521271 | 85068 | 402 | 146 | 62.8 | 298.7 |
@@ 4 Conclusion
In this article, we have demonstrated the productive combination of graph-based enumeration and neural network potentials for the computational evaluation of complex cyclization reactions with relevance to natural product synthesis planning. We implement our workflow in an open-source tool, REVAMP, which by default performs a “b4f4” and stereoenumeration scheme to generate a broad range of potential outcomes. To cope with the large number of possible elementary steps, we employ inexpensive machine learning models for intermediate filtering and perform subsequent thermodynamic and kinetic assessment with the reactive NNP AIMNet2-rxn. Many elements of this workflow take direct inspiration from YARP. We illustrate the successful construction of complex cyclization pathways with various case studies: a stereoselectivity study of an intramolecular Diels-Alder reaction and two examples from natural product syntheses representing various levels of elementary step complexity.
The summary statistics presented in Table 1 demonstrate REVAMP’s ability to explore these mechanisms with manageable computational cost. One clear trend that emerges is that the number of enumerated intermediates is very sensitive to the number of atoms which compounds the existing challenges associated with modeling larger reacting systems. REVAMP partially mitigates this combinatorial explosion with its progressive filtering stages which help reduce the number of DFT transition state optimizations by at least an order of magnitude in all case studies. Despite the efficacy of these filters, the wall time associated with performing the DFT calculations is still quite high in cases with larger molecules, highlighting the challenges associated with modeling systems of this size and complexity.
The rise of NNPs capable of providing DFT-quality energies for equilibrium and transition state geometries will continue to assist mechanistic modeling. At present, the strength of REVAMP is in the analysis of neutral, closed-shell reactions, particularly pericyclic reactions like electrocyclizations and cycloadditions. Currently, REVAMP is incapable of handling elements outside of C, H, O, and N, charged species, and open-shelled chemistry. This represents a significant limitation of its utility beyond analyzing the particular reaction types in this study. However, due to the modular nature of REVAMP, it will be straightforward to exchange the current NNP for those that will be released in coming years. In particular, we expect models trained on datasets like OMol25 42 as well as recently published AIMNet2 models, AIMNet2-Pd 85 and AIMNet2-NSE 86, to help drive the modeling of an increased reaction scope. A key consideration to ensure successful integration of these models will be ensuring these models generalize well across chemical reactions by performing activation barrier benchmarks as done in this manuscript as well as in the AIMNet2-rxn manuscript 41.
@@
The code and data that support the findings of this study are openly available at https://github.com/ncasetti7/REVAMP/tree/main. The neural network potential used in this study will be available when its corresponding paper 41 is published.